Consider the function πœ“(πœƒ) of the angular variable πœƒ, restricted to the interval βˆ’πœ‹ ≀ πœƒ ≀ πœ‹. If the wave functions satisfy the condition πœ“(πœ‹) = πœ“(βˆ’πœ‹), show that the operator 𝐿 = ℏ / 𝑖. 𝑑 / π‘‘πœƒ has a real expectation value.


1
Expert's answer
2021-09-14T09:36:08-0400

The equation for eigenfunctions is given by

"L\psi=\lambda\psi"

"\frac{\hbar}{i}\frac{d}{d\theta}\psi=\lambda\psi"

"\psi(\theta)=Ce^{i/\hbar\lambda\theta}"

"\psi(\pi)=Ce^{i/\hbar\lambda \pi},\quad \psi(-\pi)=Ce^{-i/\hbar\lambda \pi}"

"e^{i/\hbar\lambda \pi}=e^{-i/\hbar\lambda \pi}"

"e^{2i/\hbar\lambda \pi}=1"

"2/\hbar\lambda \pi=2\pi n"

The expectation values of operator

"\lambda=\hbar n=\rm real"